particle at rest decays into two particles of rest mass mo and 2mo, lighter particle moves with a speed of 0.8c, find the speed of other particle in lab frame and hence find the rest mass of the original particle

1
Expert's answer
2021-12-22T13:07:04-0500

"p_1=p_2,"

"\frac{0.8m_0c}{\sqrt{1-0.8^2}}=\frac{2m_0v_2}{\sqrt{1-\frac{v_2^2}{c^2}}},\implies""v_2=0.55c."